OpenStudy (anonymous):

Evaluate (647)(0.00916) using logarithms.

OpenStudy (anonymous):

a. 0.8490 b. 9.8490 - 10 c. 0.7728 d. 9.7728 - 10

OpenStudy (anonymous):

well it is none of these for sure

OpenStudy (anonymous):

is there a typo here? it just asks you to multiply these numbers together?

OpenStudy (anonymous):

Use logarithms to evaluate (647)(0.00916)

OpenStudy (anonymous):

charlie check that again

OpenStudy (anonymous):

Use logarithms to evaluate (647)(0.00916). The common logarithm of the answer is... a. b. c. d.

OpenStudy (anonymous):

when I take those and times them together I get 5.2652 take the log of this and that should give the answer

OpenStudy (anonymous):

what answer do you have?

OpenStudy (anonymous):

-0.824

OpenStudy (anonymous):

Oh wait, .721?

OpenStudy (anonymous):

I got it do like this take the of each number and than times them

OpenStudy (anonymous):

Take what?

OpenStudy (anonymous):

log of the first one and than take second and take those results and times them

OpenStudy (anonymous):

take thelog of 647

OpenStudy (anonymous):

what you get?

OpenStudy (anonymous):

2.81

OpenStudy (anonymous):

oh you want the log of the answer!

OpenStudy (anonymous):

now do the same for the second number

OpenStudy (anonymous):

write down the first answer

OpenStudy (anonymous):

-2.03 for the second?

OpenStudy (anonymous):

now times them togehter and take the log again

OpenStudy (anonymous):

I got -5.7043 but I got error when log

OpenStudy (anonymous):

just add, don't multiply

OpenStudy (anonymous):

\(\log(ab)=\log(a)+\log(b)\)

OpenStudy (anonymous):

yes that is right opps.

OpenStudy (anonymous):

in any case in the modern age you can do it all in one step http://www.wolframalpha.com/input/?i=log%28+%28647%29%280.00916%29+%29&a=*FunClash.log-_*Log10.Log-

OpenStudy (anonymous):

i get .7728

OpenStudy (anonymous):

but due to the () I thought it was times not add.

OpenStudy (anonymous):

satelitle73 how did come to add when there are () with numbers inside, most time that means times.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

charlie he/she is right

OpenStudy (anonymous):

Oh yes, I see answer c. matches

OpenStudy (anonymous):

sorry about all the extra the ( ) confused me.

OpenStudy (anonymous):

when you (5) (120) this means to times them.

OpenStudy (anonymous):

that is why I went that way from the start. Sorry again about all the extra work

OpenStudy (anonymous):

No, it's fine. Thank you :)

OpenStudy (anonymous):

again it's my pleasure to help as always. :)